# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v1

@@ III.3 Experimental setup and Implementation
Figure 5: Comparison of hyperparameter importances for optimizing the five-fold cross-validation score in the corresponding hyperparameter optimizations of the classification tasks of this study. For both QSVC-FQK and QSVC-PQK, the results are aggregated across all datasets within the correlating dataset family and all encoding circuits. The results corresponding to the two curves diff family are shown in (a). The hidden manifold diff dataset family is depicted in (b). In both cases we impose that $n_{\mathrm{qubits}}$ can only be integer multiples of the number of features ($d=4$), with a maximum of $n_{\text{qubits}}^{\text{max}}=15$.
IV Results
In this section, we discuss all findings from the different types of experiments that have been performed, before we summarize these results in terms of identifying general patterns in Sec. V.
IV.1 Hyperparameter Importances
[…]
Regression Tasks
@@ IV.1 Hyperparameter Importances
Finally, we note that the largest differences to other results occur for PQK-QSVR within QFMNIST. Here, the regularization $C$ shows a vanishing median value and an increased importance of both $\varepsilon$ and $w_{\mathrm{e}}$ with other parameters being similar important. In this case, the differences between hyperparameter importances on average seem to be reduced towards all parameters becoming equally important.
Classification Tasks
Figure 5 displays the hyperparameter importances for the two curves diff and hidden manifold diff dataset families. Here, for both FQK and PQK QKMs, we aggregated results across all corresponding data subsets and encoding circuits. The results in Figs. 5 (a) and (b) clearly reveal that $w_{\mathrm{e}}$ is most important in order to obtain sufficiently well trained quantum classifiers. This is in contrast to the regression results, where regularization or the $\gamma$ parameter are more important. The importance of feature scaling can be directly related to the concept of kernel bandwidth tuning as reported, e.g., in Refs. [26, 27] and is in accordance with the results presented therein. Moreover, it perfectly agrees with the recent outcomes of Ref. [37], where the scaling parameter $t$ of the Hamiltonian evolution feature map was evaluated as the most important parameter in PQKs for both the accuracy as well as for the geometric difference. In addition to that, we provide some evidence that the importance of feature prescaling is slightly more pronounced for FQKs than for PQKs. This, however, is not surprising, given the fact that the $\gamma$-parameter in PQKs of the form of Eq. (9) acts as an additional (classical) bandwidth tuning parameter. Here, $\gamma$ is the second most important parameter for PQK models. The regularization parameter $C$ of QSVC appears to be comparatively unimportant, except for some outliers. This is even more pronounced for PQKs than for FQKs. Compared to Fig. 4, the results in Fig. 5 clearly indicate that $n_{\text{qubits}}$ and $n_{\text{layers}}$ appears to be more important, with tuning the number of layers being more important than potentially encoding features redundantly on the number of qubits. This could hint at the more difficult the dataset, which for $D,m\geq 6$ is the case for the classification datasets (cf. Fig. 3), the more important feature redundancies become and adjusting the number of layers accordingly.
IV.2 Model Performance and Correlation Analysis
Figure 6: Overview of test performance scores of respective QKMs as a function of increasing dataset complexity. Results within each dataset are aggregated across all data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$ yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays two regression tasks, where the MSE is used to measure the perdiction accuracy. The Friedman dataset family is shown in (a), where the number of features $n$ can be viewed as control parameter for dataset complexity. The QFMNIST dataset family is shown in (b), where the number of principal components used in the dataset generation procedure controls the dataset complexity. The lower panel illustrates the two classification tasks of this study, where we use the ROC-AUC score to assess classification accuracy. In (c) we show the two curves diff dataset family with the degree $D$ controlling the complexity. The hidden manifold diff family is given in (d) with the manifold dimension $m$ as respective control parameter.
In the preceding section, we analyzed the impact of various QKM hyperparameters on effective model tuning, specifically exploring the optimization question of how to obtain an optimally trained model. In the following, we delve into the relationships concerning the generalizability of QKMs. Due to the insights from the hyperparameter importances in Fig. 4 and 5, we fix $n_{\mathrm{qubits}}$ for every dataset to the respective number of features for all upcoming investigations and perform hyperparameter optimizations for every data encoding circuit with $n_{\mathrm{layers}}\in[1,8]$
First, we examine the test performance of different QKM models as a function of increasing dataset complexity. This analysis is given in Fig. 6, which illustrates the results for the Friedman, QFMNIST, two curves diff, and hidden manifold diff datasets. In these cases results for each QKM and dataset within the respective dataset family are aggregated across the data encoding circuits with corresponding obtained optimal $n_{\mathrm{layers}}^{*}$, yielding the best MSE or ROC-AUC score, respectively.
A comparison with Fig. 3 indicates that the behavior of test performance scores as a function of a control parameter qualitatively aligns with the increasing complexity of the respective datasets. This confirms that $\overline{C}$ is a useful measure of complexity. The comparison between FQK and PQK approaches in Figs. 6 (a) and (b), particularly for larger problem instances with increasing $n$, which transfers to increasing $n_{\mathrm{qubits}}$, reveals no significant performance differences. While PQKs demonstrate slightly superior performance, the difference is not as substantial as one might anticipate given the challenges posed in connection with exponential concentration [28], whose regime we may already enter with up to 15 qubits. The absence of this observation may be explained by the presence of the bandwidth tuning parameter $w_{\mathrm{e}}$, which shows significant correlations with performance scores as discussed below. The comparison between QKRR and QSVR each within FQK or PQK approaches does not reveal a clear advantage for either method. For the Friedman dataset we notice comparatively large interquartile ranges (IQRs) for QKRR/QSVR-PQK approaches towards datasets with lower dataset complexity, especially for $n=6$. By inspecting the corresponding training scores in Fig. A10 (a) in Appendix D.1, we may attribute this trend to poorly trained models for QSVR-PQK, while for QKRR-PQK this may be due to overfitting. The latter could probably be alleviated by increasing the number of training data which is beyond the scope of this work. Moreover, the QFMNIST($n=2$) dataset constitutes the only case for which we observe a clear performance difference between FQK and PQK. Regarding the classification tasks, we do not realize any significant difference between both QKMs.
Beyond that, we generally observe many outliers exhibiting particularly poor test performance scores. The comparison with associated training performance scores in Fig. A10 shows that this may again result from overfitting for certain model combinations which might be too expressive for the respective dataset. This may also suggest that the design of data encoding circuits should follow a data-centric approach as, e.g., pioneered in Ref. [24]. Finally, we note that the propensity for overfitting is not only present for the two curves diff, as previously noted in Ref. [34], we also observe this for the Friedman and QFMNIST regression families. As shown in Appendix D.1 this is most dominant for QKRR-FQK/PQK approachs with increasing dataset complexity, while QSVR appears to be more robust. The hidden manifold diff dataset, on the other hand, is not as prone to overfitting. However, in total we point out here that our test size is quite small ($M^{\prime}=60$, as mentioned in Sec. III.2) which represents a clear limitation towards more detailed statements in this direction. This should be carefully considered in future studies.
To investigate the impact of data encoding circuits on the performance we aggregate results across all datasets within each family and $n_{\mathrm{layers}}\in[1,8]$ and then re-plot the data, ordered by the encoding circuits. This is exemplified in Fig. 7 for the Friedman (regression) and the two curves diff dataset (classification). The results corresponding to the other dataset families of this study are shown in Fig. A11 in Appendix B. When examining the performance scores for a single dataset within a dataset family, significant performance differences emerge. This can be seen in Figs. 7 (a) and (c). Here, the “HZY_CZ_EncodingCircuit” performs particularly poor in regression tasks, while the “ChebyshevPQC encoding circuit” under performs in classification tasks, both for FQK and PQK approaches. However, when considering performance scores aggregated across all datasets within a dataset family, we observe that, except for the “ChebyshevPQC” in the QSVC-FQK case, all encoding circuits perform comparably, cf. Figs. 7 (b) and (d). No clear winner emerges, and in fact, each encoding circuit has upper whiskers indicating perfect test performance scores. The IQRs are also very similar across all encoding circuits. Furthermore, this aggregated analysis does not reveal any significant performance differences between FQK and PQK approaches. The similarity in the aggregated test performance scores is more pronounced in the classification problem illustrated (see Appendix D for analogous plots for other datasets). This again supports the argument for problem-specific data encoding architecture search [43, 24].
Figure 7: Insight into the impact of data encoding circuits on test performance scores. Results are aggregated across all datasets within the respective family and $n_{\mathrm{layers}}\in[1,8]$. The upper panel exemplifies this for the Friedman regression problem. In (a) we show results corresponding to the dataset with $n=15$ features. Here, test performance scores, as measured by MSE, are aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel (FQK/PQK). The findings in (b) represent test performance scores aggregated over all datasets in the Friedman family. The lower panel details the two curves diff dataset. In (c), we display test performances, as measured by ROC-AUC score, aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with degree $D=13$. Aggregated results across all datasets are depicted in (d).
In earlier studies [34], it has been found that data encoding circuits without entanglement can perform surprisingly well. In our case, these are the “SeparableRx” and “ZFeatureMap” encoding circuits. As demonstrated in Ref. [27] for FQKs, the “SeparableRx” circuit is analytically tractable, and this can be straightforwardly extended to PQKs. Similarly, the “ZFeatureMap”, which incorporates additional Hadamard gates, also demonstrates comparable performance. This raises the possibility that driving forces other than “quantumness” might be fostering the performance of QKMs. This observation underscores the need for further comprehensive and systematic investigations into the underlying factors contributing to QKM performance.
To gain even deeper insights, Fig. 8 provides an exemplary overview on the results of the correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves for the QFMNIST dataset family. We refer to Fig. A12 in the Appendix D.1 for a summary of all other dataset families investigated in this study. The results are aggregated across datasets within a dataset family and across all considered encoding circuits and $n_{\mathrm{layers}}\in[1,8]$, respectively. Due to the presence of numerous outliers, cf. Figs. 6 and 7, we employ Spearman correlation analysis, which does not assume normally distributed variables or that the relationship between variables is linear. Figure 8 indicates statistically significant correlations by green-highlighted $p$-values. Note that the signs of the correlation coefficients work in opposite directions for regression and classification tasks, due to the use of MSE and ROC-AUC metrics, respectively.
Figure 8: Overview of the Spearman correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves, for the QFMNIST dataset family. The results correspond to aggregating data across all datasets, encoding circuits, and $n_{\mathrm{layers}}\in[1,8]$. The Spearman correlation coefficients are given on the lower triangles of the respective matrices (with blue for negative to red for positive coefficients), while the upper triangles display corresponding $p$-values. Here, statistically significat correlations ($p\leq 0.05$) are highlighted in green. Since we use the MSE to evaluate regression performance and the ROC-AUC score in case of classification, respectively, we note that the signs of associated correlation coefficients work in opposite direction; i.e., e.g., negative correlation with MSE means better regression performance, while it is the other way round for classification.
In line with the findings from the analysis of hyperparameter importance in Figs. 4 and 5, we observe moderate to strong statistically significant correlations between test performance scores and the various regularization parameters, i.e., $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively as well as $\varepsilon$ for QSVR. Additionally, consistent with findings on bandwidth tuning [26, 27], we frequently (also considering the findings from Fig. A12) observe moderate and statistically significant correlations between the feature scaling parameter $w_{\mathrm{e}}$, cf. Eq. (10), and and performance scores. This parameter often also correlates with the $n_{\mathrm{layers}}$. For QSVR-/QSVC-PQK models, moderate correlations are also observed between the length-scale parameter $\gamma$ of the Gaussian outer kernel and test performance. With the complete correlation analyses in mind, cf. Fig. A12, only in the case of the artificial quantum dataset QFMNIST, we observe very weak correlations between $n_{\mathrm{layers}}$ and the test MSE.
Given these insights, we may acquire an improved understanding of the poor performance regimes observed in Figs. 6 and 7 if we interpret the strong correlations between regularization parameters and model performance as indicative of ill-conditioned training kernel Gram matrices. This could be a result of unfavorable embeddings of the underlying data into the quantum Hilbert space. Combined with the obtained correlations between performance scores and $w_{\mathrm{e}}$ as well as $\gamma$, this could further challenge our understanding of “quantumness” of QKMs, apart from the already mentioned entanglement problem.
IV.3 Data Encoding Strategies
It is known [11, 14] that the number of qubits and layers of data encoding circuits determine the Fourier frequency spectrum of the resulting quantum model and that encoding features redundantly can help in getting more accurate results. However, there are some degrees of freedom how to do this exactly. Therefore, we try shed some light on this in the context of QKMs. To this end, we study quantum kernel models built from encoding circuits according to the two data embedding schemes illustrated in Fig. 9. In both embedding schemes, features are encoded in the qubits from top to bottom. Differences in the two encoding strategies occur when the number of qubits is larger than the number of features. In “option 1”, after all features have been encoded once in a layer, the encoding is repeated and then cut off once the final qubit is reached. In “option 2”, the features enumeration of features is not reset at the next layer. As a consequence, each feature in “option 1” is always assigned to the same qubit, whereas in “option 2” the features can be shuffled across qubits.
Figure 9: Schematic illustration of two different data encoding options considered in the investigations of Sec. IV.3.
In the case of the regression (QKRR and QSVR), Fig. (10) shows exemplary the result for the QFMNIST dataset for $n=5$. The results for each quantum kernel were aggregated across the regression methods and the respective encoding circuits. Then, the median of the resulting MSE test score was calculated. For FQK, both embedding options show that largest scores are achieved for ten qubits, i.e., for integer multiples of the number of features. It can be seen that higher scores occur more frequently with embedding “option 2”. For PQK, it appears that no feature redundancies are necessary in either case. Here, however, embedding “option 1” often delivers larger scores.
Figure 10: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the QFMNIST regression dataset with $n=5$ principal components. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
Figure 11 displays the results for the two curves diff classification dataset for $D=13$ (the number of features is still fixed to $d=4$). The results were aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores was evaluated. It can be seen that here, unlike in the case of regression, both embedding options for both quantum kernels (FQK and PQK) deliver the best classification accuracies for non-integer multiples of the number of features. Here, “option 1” performs better in each case. In addition, there seems to be a tendency for the larger $n_{\text{layers}}$ and the larger $n_{\text{qubits}}$ the better the classification accuracy. This result might be due to the much greater dataset complexity the two curves diff ($D=13$) dataset compared to QFMNIST with $n=5$.
Figure 11: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK- and PQK-QSVC approaches, respectively for the two curves diff classification dataset with $D=13$. Note, the feature dimension is kept constant at $d=4$. The results were aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
However, the findings discussed here also apply to the remaining regression and classification datasets in this study. These are summarized in detail in Appendix D.2 in Figs. A13A14 and A15.
IV.4 In-depth Analysis of PQK Design Options
Beyond tuning various hyperparameters and selecting appropriate data encoding strategies, QKMs based on PQKs according to Eq. (8) have additional degrees of freedom in model design. Specifically, any outer kernel function $\kappa$ that satisfies the requirements of a kernel function can be applied to the features computed by quantum inputs according to Eq. (II). Here, the exact form of the measurement operator $O^{k}$ used to define the projected quantum circuits is not fixed either. To the best of our knowledge, a thorough investigation of these aspects, aside from Ref. [37], which considers the impact of the RDM size $k$ on cross-validation accuracy and generalization ability, is still lacking.

### v2

@@ IV.2 Influence of Hyperparameters
Figure 6: Comparison of hyperparameter importances for optimizing the five-fold cross-validation score in the corresponding hyperparameter searches for the classification tasks of this study. The results for each model are aggregated across all datasets within the correlating dataset family and all encoding circuits. The results corresponding to the two curves diff family are shown in (a). The hidden manifold diff dataset family is depicted in (b). In both cases we impose that $n_{\mathrm{qubits}}$ can only be integer multiples of the number of features ($d=4$), with a maximum of $n_{\text{qubits}}^{\text{max}}=15$.
Figure 7: Overview of the Spearman correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves, for the QFMNIST dataset family. The results correspond to aggregating across all datasets, encoding circuits, and $n_{\mathrm{layers}}\in[1,8]$. The Spearman correlation coefficients are given on the lower triangles of the respective matrices (with blue for negative to red for positive coefficients), while the upper triangles display corresponding $p$-values. Here, statistically significat correlations ($p\leq 0.05$) are highlighted in green.
@@
For classifications tasks, Fig. 6 displays the hyperparameter importances for the two curves diff and hidden manifold diff dataset families. Here, for both FQK and PQK QKMs, we aggregate results across all corresponding data subsets and encoding circuits. The results in Figs. 6 (a) and (b) clearly reveal that $w_{\mathrm{e}}$ is most important in order to obtain well trained quantum classifiers, which is in contrast to the regression results. Again, the importance of feature scaling can be directly related to the concept of kernel bandwidth tuning [26, 27] and is in accordance with the results presented therein. Moreover, it perfectly agrees with the recent outcomes of Ref. [37], where the scaling parameter $t$ of the Hamiltonian evolution feature map was evaluated as the most important parameter in PQKs for both the accuracy as well as for the geometric difference of selected classification tasks. In addition to that, we provide some evidence that the importance of feature prescaling is slightly more pronounced for FQKs than for PQKs. This, however, is not surprising, given the fact that the $\gamma$-parameter in PQKs of the form of Eq. (9) acts as an additional (classical) bandwidth tuning parameter. Here, $\gamma$ is the second most important parameter for PQK models. The regularization parameter $C$ of QSVC appears to be comparatively unimportant, except for some outliers. This is even more pronounced for PQKs than for FQKs. In comparison to Fig. 5, the results in Fig. 6 clearly indicate that $n_{\text{qubits}}$ and $n_{\text{layers}}$ appear to be more important, with tuning the number of layers being slightly more important than potentially encoding features redundantly on the number of qubits. This may be due to the increased dataset complexity for $D,m\geq 6$, cf. Fig. 3, which might require more expressive models. This can be achieved by introducing feature redundancies and adjusting the number of layers accordingly [11, 14].
Nevertheless, given the relatively low median values for the hyperparameter importance of $n_{\mathrm{qubits}}$ in both Figs. 5 and 6, we a posteriori justify our previous choice of fixing this parameter in the respective investigations of model performance.
To obtain even deeper insights into the influence of hyperparameters, Fig. 7 provides an exemplary overview on the results of the correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves for the QFMNIST dataset family. We refer to Fig. A11 in Appendix D.2 for a summary of all other dataset families investigated in this study. The results are aggregated across datasets within a dataset family and across all nine encoding circuits and $n_{\mathrm{layers}}\in[1,8]$, respectively. As a measure for correlation, we calculate the Spearman coefficient, which neither requires normally distributed variables nor a linear relationship between them and is more robust to outliers. Figure 7 indicates statistically significant correlations by green-highlighted $p$-values. In line with the findings from the analysis of the importance of hyperparameter in Figs. 5 and 6, we observe moderate to strong statistically significant correlations between test performance scores and the various regularization parameters, i.e., $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively as well as $\varepsilon$ for QSVR. Additionally, consistent with findings on bandwidth tuning [26, 27], we frequently observe moderate and statistically significant correlations between the feature scaling parameter $w_{\mathrm{e}}$, cf. Eq. (10), and performance scores (also considering the findings from Appendix D.2 in Fig. A11). This parameter often also correlates with the $n_{\mathrm{layers}}$. For QSVR-/QSVC-PQK models, moderate correlations are also observed between the length-scale parameter $\gamma$ of the Gaussian outer kernel and test performance. Finally, we note that the $n_{\mathrm{layers}}$ parameter only shows weak statistically significant correlations to the test performance for the QFMNIST dataset.
As mentioned above, the correlation analysis provides an explanation for the poor performance regimes observed in Fig. 4. We interpret the strong correlations between regularization parameters and model performance as indicative of ill-conditioned training kernel Gram matrices. This could be a result of unfavorable embeddings of the underlying data into the quantum Hilbert space. Furthermore, it corroborates the observed tendency to overfitting of some models.
IV.3 Influence of Encoding Circuits
Figure 8: Insight into the impact of data encoding circuits on test performance scores. Results are aggregated across all datasets within the respective family and $n_{\mathrm{layers}}\in[1,8]$. The upper panel exemplifies this for the Friedman regression problem. In (a) we show results corresponding to the dataset with $d=15$ features. Here, test performance scores, as measured by MSE, are aggregated over $n_{\mathrm{layers}}\in[1,8]$ per regression method (QSVR/QKRR) and quantum kernel (FQK/PQK). The findings in (b) represent test performance scores aggregated over all datasets in the Friedman family. The lower panel details the two curves diff dataset. In (c), we display test performances, as measured by ROC-AUC score, aggregated over $n_{\mathrm{layers}}\in[1,8]$ for the dataset with degree $D=13$. Aggregated results across all datasets are depicted in (d).
To investigate the impact of data encoding circuits on the performance of the resulting models, we aggregate results across all datasets within each family and choose $n_{\mathrm{layers}}$ varying from 1–8. We then show the results for each encoding circuit. This is exemplified in Fig. 8 for the Friedman (regression) and the two curves diff (classification) datasets. The results corresponding to the other dataset families of this study are shown in Fig. A12 of the Appendix D.3. When examining the performance scores for a single dataset within a dataset family, significant performance differences among the data encoding circuits emerge. This can be seen in Figs. 8 (a) and (c). Here, the “HighDimEncodingCircuit” and the “ZZFeatureMap” perform particularly poor in regression tasks, while the “ChebyshevPQC encoding circuit” underperforms in classification tasks, both for FQK and PQK approaches (note that trends for MSE and ROC-AUC performance scores work in opposite directions). However, when considering performance scores aggregated across all datasets within a dataset family, we observe that, except for the “ChebyshevPQC” in the QSVC-FQK case and with some trend for the “ZZFeatureMap” in regression tasks, all encoding circuits perform comparably in terms of overall performance and variance (cf. Figs. 8 (b) and (d)). This supports arguments for problem-specific data encoding architecture search [43, 24]. Furthermore, this aggregated analysis does not reveal any significant performance differences between FQK and PQK approaches.
In earlier studies [34], it has been found that data encoding circuits without entanglement can perform surprisingly well. In our case, these are the “SeparableRx” and “ZFeatureMap” encoding circuits. As demonstrated in Ref. [27] for FQKs, the “SeparableRx” circuit is analytically tractable, and this can be straightforwardly extended to PQKs. Similarly, the “ZFeatureMap”, which incorporates additional Hadamard gates, also demonstrates comparable performance. With the model performances obtained in Fig. 8, this raises the possibility that driving forces other than “quantumness” might be fostering the performance of QKMs. This observation underscores the need for further comprehensive and systematic investigations into the underlying factors contributing to QKM performance.
IV.4 Influence of Data Encoding Strategies
As already implicitly indicated above, it is known that the number of qubits and layers of data encoding circuits determine the Fourier frequency spectrum of the resulting quantum model and that encoding features redundantly can help in getting more accurate results [11, 14]. However, from a practical point of view it is not clear how to best distribute the features of a data point among the available qubits in a particular encoding circuit. To shed some light on this aspect, we compare QKMs built from selected encoding circuits following the two encoding schemes illustrated in Fig. 9. In both embedding schemes, features are encoded in the qubits from top to bottom. Differences in the two encoding strategies occur when the number of qubits is larger than the number of features. In “option 1”, after all features have been encoded once in a layer, the encoding is repeated and then cut off once the final qubit is reached. In “option 2”, the enumeration of features is not reset at the next layer. As a consequence, each feature in “option 1” is always assigned to the same qubit, whereas in “option 2” the features can be shuffled across qubits.
Figure 9: Schematic illustration of two different data encoding options considered in the investigations of Figs. 10 and  11.
Using the QFMNIST dataset with $d=5$ as an example, Fig. (10) shows the findings for this regression tasks. The results for each quantum kernel are aggregated across the regression methods and the respective encoding circuits. Then, the median of the resulting MSE test score is calculated. For FQK, both embedding options show that largest scores are achieved for ten qubits, i.e., for integer multiples of the number of features. It can be seen that higher scores occur more frequently with embedding “option 2”. For PQK, it appears that no feature redundancies are necessary in either case. Here, however, embedding “option 1” often delivers larger scores. Overall, the MSE differences are small.
Figure 10: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the QFMNIST regression dataset with $d=5$ principal components. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
Figure 11 displays the results for the two curves diff classification dataset for $D=13$ (the number of features is still fixed to $d=4$). Again, results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is evaluated. It can be seen that here in contrast to the previous example, both embedding options deliver the best classification scores for non-integer multiples of the number of features (for PQK and FQK). Here, “option 1” performs better in each case. In addition, larger values of $n_{\text{layers}}$ and $n_{\text{qubits}}$ lead to increasing classification scores. The difference to the previous example (Fig. 10) might be due to significantly larger dataset complexity. This also explains the overall larger difference in performance scores for the two curves diff .
Figure 11: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK- and PQK-QSVC approaches, respectively for the two curves diff classification dataset with $D=13$. Note, the feature dimension is kept constant at $d=4$. The results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
The trends discussed here also apply to the other datasets in this study. These are summarized in detail in Appendix D.4.
IV.5 Analysis of PQK Design Options
Projected quantum kernels as defined in Eq. (8) have additional degrees of freedom which can strongly influence the performance of the resulting model. Specifically, any proper outer kernel function $\kappa$ can be applied and the measurement operator $O^{k}$ can be chosen freely. To the best of our knowledge, a thorough investigation of the influence of both of these design choices on the model is still lacking. [^4] Reference [37] touches upon similar aspects but has a much narrower focus on the impact of the RDM size on cross-validation accuracy and generalization ability.
[…]